# On a class of toric manifolds arising from simplicial complexes

### v1

@@ Abstract.
Given an arbitrary abstract simplicial complex $K$ on $[m]:=\{1,2,\ldots,m\}$, different from the simplex $\Delta_{[m]}$ with $m$ vertices, we introduce and study a canonical $(2m-2)$-dimensional toric manifold $X_{K}$, associated to the canonical $(m-1)$-dimensional complete regular fan $\Sigma_{K}$. This construction yields an infinite family of toric manifolds that are not quasitoric and provides a topological proof of the Dehn-Sommerville relations for the associated Bier sphere $\mathrm{Bier}(K)$. Finally, we prove a criterion for orientability of canonical real toric manifolds.
@@ 1. Introduction
$(\underline{X},\underline{A})^{K}=\mathop{\bigcup}_{I\in K}\{(x_{1},\ldots,x_{m})\,|\,x_{i}\in A_{i}\text{ if }i\notin I\}\subseteq\prod\limits_{i=1}^{m}X_{i},\text{ where }\underline{X}=\{X_{i}\,|\,1\leq i\leq m\},\,\underline{A}=\{A_{i}\,|\,1\leq i\leq m\},$
and $(X_{i},A_{i})$ is a topological pair for each $1\leq i\leq m$, and combinatorial properties of $K$ is one of the key problems considered in toric topology [6]. In the particularly important and well-studied cases of a moment-angle complex $\mathcal{Z}_{K}=(\mathbb{D}^{2},\mathbb{S}^{1})^{K}$ and its real analogue $\mathcal{R}_{K}=(\mathbb{D}^{1},\mathbb{S}^{0})^{K}$, see [4], it turns out that these polyhedral products are not merely cellular spaces, but also acquire a structure of a topological, or even a smooth manifold consistent with the coordinate torus action, provided $K$ is a sphere, see Definition 2.3 and Theorem 4.7.
Moreover, if the Buchstaber number $s(K)$ of a starshaped sphere $K$ (i.e. the maximal rank of a toric subgroup $H$ in the coordinate torus $\mathbb{T}^{m}$ acting freely on $\mathcal{Z}_{K}$) is maximal possible (i.e. equals $m-n$, where $n=\dim K+1$), then the topological toric manifold $\mathcal{Z}_{K}/H$ is defined. This smooth $2n$-dimensional manifold with an $n$-dimensional compact torus $T^{m}/H$ action is a topological counterpart of a toric manifold: a complete nonsingular toric variety $X_{\Sigma}$ of complex dimension $n$, which is defined in the framework of toric geometry [10] in terms of a complete regular fan $\Sigma$ in ${\mathbb{R}}^{n}$, see Definition 2.8. The real analogues of the Buchstaber invariant and the (topological) toric manifold are defined similarly.
@@
By definition, first introduced in the unpublished note [2], given an arbitrary simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$, its Bier sphere $\mathrm{Bier}(K)$ equals the deleted join of the complex $K$ and its Alexander dual complex $K^{\vee}$, see Definition 2.4. Different proofs of the fact that this construction yields a PL-sphere of dimension $(m-2)$ appeared in [21] and [20]. More recently, a canonical starshaped realization for any Bier sphere was obtained in [15], see also [16] and Theorem 2.6. On the other hand, it was observed in [21] that almost all Bier spheres of simplicial complexes on $[m]$ are non-polytopal, as $m\to\infty$; however, no particular example of a non-polytopal Bier sphere has been constructed so far.
@@
In this paper, we begin with a canonical regular realization for any Bier sphere and discuss some applications of our construction in topology and combinatorics. This allows us to generalize (refine) the results on the canonical starshaped realization [15, 16] and the Buchstaber numbers computation for Bier spheres [17, 18].
The structure of the paper is as follows. In Section 2, we collect all the necessary definitions, constructions, and results related to simplicial complexes and fans. Based on that, we introduce the notion of a Bier sphere and recall its basic combinatorial and geometric properties. The central result in Section 3 is Theorem 3.3. Here we focus on the regularity of the canonical, complete fan $\Sigma_{K}$ (the radial fan of the starshaped realization of $\mathrm{Bier}(K)$, described in Theorem 2.6). As a first application we observe that whenever $\mathrm{Bier}(K)$ is polytopal the combinatorial simple polytope $P$ such that $\mathrm{Bier}(K)\cong\partial P^{*}$ is Delzant. In Section 4, we introduce the canonical (real) moment-angle manifolds and (real) toric manifolds corresponding to the canonical fan $\Sigma_{K}$ and discuss further applications of our main construction. First, we obtain a topological proof of the Dehn-Sommerville relations for Bier spheres based on Poincaré duality for toric manifolds (Proposition 4.3). Then an infinite family of toric manifolds that are not quasitoric is identified taking into account the canonical toric manifolds over non-polytopal Bier spheres (Proposition 4.5). Finally, a criterion for orientability of the canonical real toric manifolds over Bier spheres is proved based on the canonical characteristic matrix $\Lambda_{K}$ construction (Proposition 4.11).
@@ Acknowledgements
The authors are grateful to Anton Ayzenberg, Taras Panov, and Aleš Vavpetič for various fruitful discussions, valuable comments and suggestions. The authors were supported by the Serbian Ministry of Science, Innovations and Technological Development through the Mathematical Institute of the Serbian Academy of Sciences and Arts.
@@ 2. Abstract simplicial complexes and rational polyhedral fans
Set $[m]:=\{1,2,\ldots,m\}$ for an integer $m\geq 1$. We say that a non-empty subset $K\subseteq 2^{[m]}$ is an (abstract) simplicial complex on $[m]$ if $\sigma\in K$ and $\tau\subseteq\sigma$ imply $\tau\in K$. Elements of $K$ are called its simplices, or faces. Faces of cardinality $1$ are called (geometric) vertices. The set of vertices of $K$ will be denoted by $V(K)$.
@@
Given two simplicial complexes, $K_{1}$ and $K_{2}$, on $[m]$, we say that they are isomorphic and write $K_{1}\cong K_{2}$, if there exists a bijection $\phi\colon[m]\to[m]$ such that both $\phi$ and its inverse $\phi^{-1}$ map simplices to simplices. For an $(n-1)$-dimensional simplicial complex $K$, we define its $f$-vector to be the $(n+1)$-tuple $(f_{-1},f_{0},\ldots,f_{n-1})$ such that $f_{i}$ equals the number of $i$-dimensional faces of $K$ for $i\geq-1$. Note that $f_{-1}=1$. Then the $h$-vector of $K$ is defined to be the $(n+1)$-tuple $(h_{0},h_{1},\ldots,h_{n})$ determined from the identity:
@@
We call a pure $(n-1)$-dimensional simplicial complex $K$
@@
Now let us introduce the main object of study in this paper, a Bier sphere. Consider a copy $[m^{\prime}]:=\{1^{\prime},2^{\prime},\ldots,m^{\prime}\}$ of the set $[m]$ with $m\geq 2$.
@@
  • the Alexander dual of $K$ is a simplicial complex $K^{\vee}$ on $[m^{\prime}]$ such that
@@
  • the Bier sphere of $K$ is a simplicial complex $\mathrm{Bier}(K)$ on the disjoint union $[m]\sqcup[m^{\prime}]$ such that
@@
Due to [17, Corollary 2.7], the Bier sphere $\mathrm{Bier}(K)$ has $m-f_{0}(K)+f_{m-2}(K)$ ghost vertices. Hence it has $f_{0}(\mathrm{Bier}(K))=2m-(m-f_{0}(K)+f_{m-2}(K))=m+f_{0}(K)-f_{m-2}(K)$ geometric vertices.
@@
It follows easily from the above definition that in dimension 1 the only Bier spheres are the boundaries of a $p$-gon with $p=3,4,5$, and $6$, see [19, Example 2.8]. In dimension 2, there exist $13$ different combinatorial types of Bier spheres, see [17, Theorem 2.16].
It was shown in [2] that $\mathrm{Bier}(K)$ defined above is a PL-sphere of dimension $(m-2)$ with the number of vertices varying between $m$ and $2m$. An explicit construction of the required PL-homeomorhism between $\mathrm{Bier}(K)$ and $\partial\Delta^{m-1}$ can be found in [20]. Here is a generalization of this result, which is crucial for this paper.
@@ 3. Regular realizations for Bier spheres
Note that the statement is obvious for $m=2$. Therefore, assume $m\geq 3$. First let us show that $\Sigma_{K}$ is a fan. By definition, $C(I,J)$ is a rational polyhedral cone generated by the set $G(I,J)$ of vectors in ${\mathbb{R}}^{m-1}$. Moreover, $G(I,J)$ is a minimal set of generators for $C(I,J)$, since $G(I,J)$ is a linearly independent set of vectors in ${\mathbb{R}}^{m-1}$.
@@
where $C^{\circ}(I,J)$ is the ”interior” of $C(I,J)\in\Sigma_{K}$, consisting of linear combinations of $-e_{i}$ and $e_{j}$ with strictly positive scalars. This would also imply that the resulting fan $\Sigma_{K}$ is complete.
Here we use a relation between Bier spheres which are sufficiently close. It is well known ([21], proof of Theorem 5.6.2) that if $K$ and $K\cup\{A\}$ are simplicial complexes with vertices in $[m]$, then:
$\mathrm{Bier}(K\cup A)=\mathrm{Bier}(K)\setminus D_{1}\cup D_{2}$
@@
    $\Sigma_{\varnothing_{[m]}}$ satisfies (3.1) (which is shown in Example 3.2 because $\Sigma_{\varnothing_{[m]}}$ is a complete regular fan);
@@
(3.2) $x=\alpha_{i_{1}}e_{i_{1}}+\cdots+\alpha_{i_{k}}e_{i_{k}}+\beta_{j_{1}}e_{j_{1}}+\cdots+\beta_{j_{l}}e_{j_{l}}$
@@
Without loss of generality we may assume that
@@
and here $\alpha_{i_{1}}-\beta_{j_{1}}\leq 0,\ldots\alpha_{i_{k}}-\beta_{j_{1}}\leq 0,-\beta_{j_{1}}\leq 0$ and $(\beta_{j_{2}}-\beta_{j_{1}})\geq 0,\beta_{j_{l}}-\beta_{j_{1}}\geq 0$ implying that $C(A,A^{c}\setminus\{j_{1}\})$ is the cone of $\mathbb{D}_{2}$ containing $x$. Because of the uniqueness of (3.3), point $x$ will be contained in two cones $C(A,A^{c}\setminus\{j_{1}\})$ and $C(A,A^{c}\setminus\{j_{2}\})$ if and only if $\beta_{j_{1}}=\beta_{j_{2}}$ so by disregarding the coefficients which are equal to zero in (3.3), we obtain the unique cone of $\Sigma_{K\cup\{A\}}$ containing $x$ in its interior.
@@
Since $-\beta_{j_{1}}\leq 0,\alpha_{i_{2}}-\beta_{j_{1}}\leq 0,\ldots,\alpha_{i_{k-1}}-\beta_{j_{1}}\leq 0,\alpha_{m}-\beta_{j_{1}}\leq 0$ and $(\beta_{j_{2}}-\beta_{j_{1}})\geq 0,\ldots,\beta_{j_{l}}-\beta_{j_{1}}\geq 0$, we conclude that $C(A,A^{c}\setminus\{j_{1}\})$ is the cone of $\mathbb{D}_{2}$ containing $x$. The uniqueness of the cone of $\Sigma_{K\cup\{A\}}$ containing $x$ in its interior follows as in the previous case.
@@
proving that the cone $C(A,A^{c}\setminus\{m\})$, and therefore its corresponding sub-cone, contains point $x$ in its interior.
@@
As an immediate corollary we obtain the following statement.
@@ 4. Applications
Observe that a nonsingular projective toric variety $X_{\Sigma}$ corresponding to the normal fan $\Sigma=\Sigma_{P}$ of an $n$-dimensional Delzant polytope $P$ is a quasitoric manifold of real dimension $2n$ over the combinatorial polytope $P$ via restricting the action of $(\mathbb{C}^{*})^{n}$ to $\mathbb{T}^{n}\subset(\mathbb{C}^{*})^{n}$. In what follows we consider toric manifolds with this restricted action and we denote them by $M_{\Sigma}$. The existence of the class of quasitoric manifolds arising as toric manifolds over Delzant polytopes motivated the problem, see [5, Problem 6.26], of constructing a toric manifold, which is not a quasitoric manifold. In [25], it was done by a consecutive resolution of singularities for a toric variety over the Barnette nonpolytopal 3-sphere with 8 vertices. The canonical complete regular fan $\Sigma_{K}$ construction yields a different approach to this problem.
@@
There are infinitely many toric manifolds whose orbit spaces are not homeomorphic to any simple polytope as manifolds with corners.
@@
  • the smooth $(m-1)$-dimensional canonical real moment-angle manifold $\mathcal{R}_{\mathrm{Bier}(K)}$ over $K$.
@@
In the above definition, we apply Theorem 4.7 when $m\geq 4$. If $m\leq 3$, then $\mathrm{Bier}(K)$ is a polytopal sphere, and therefore its (real) moment-angle complex acquires an equivariant smooth structure due to [7].
@@
$i\mapsto-e_{i}\text{ for all }1\leq i\leq m\text{ such that }i\in V(\mathrm{Bier}(K)),\text{ and }i^{\prime}\mapsto e_{i}\text{ for all }1\leq i\leq m\text{ such that }i^{\prime}\in V(\mathrm{Bier}(K)).$
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Define the canonical (real) toric manifold over $K$ as a quotient space of the canonical (real) moment-angle manifold over $K$:
@@
Due to [22, Theorem 1.7], the small cover $M_{{\mathbb{R}}}(P,\bar{\Lambda}_{K})$ is orientable if and only if the image of the composition $\epsilon\bar{\Lambda}_{K}$ equals $\{1\}$. This implies the statement, since ±$\pm e_{m}=\mp(e_{1}+\cdots+e_{m-1})\in{\mathbb{R}}^{m-1}$. If $\Sigma_{K}$ is not a normal fan of a Delzant polytope, a straightforward modification of the proof of [22, Theorem 1.7] with the starshaped region $P$ bounded by the canonical starshaped realization of $\mathrm{Bier}(K)$ constructed in [15] shows the conclusion of the orientability criterion is still valid in our case, which finishes the proof. ∎

### v2

@@ Abstract.
Given an arbitrary abstract simplicial complex $K$ on $m]:=\{1,2,\ldots,m\}$, different from the simplex $\Delta_{[m]}$ with $m$ vertices, we introduce and study a canonical (2m2)(2m-2)-dimensional toric manifold $X_{K}$, associated to the canonical (m1)(m-1)-dimensional complete regular fan $\Sigma_{K}$. This construction yields an infinite family of toric manifolds that are not quasitoric and provides a topological proof of the Dehn-Sommerville relations for the associated Bier sphere $\mathrm{Bier}(K)$. Finally, we classify the canonical real and complex moment-angle manifolds of Lusternik-Schnirelmann category $\leq 2$ and prove a criterion for orientability of the canonical real toric manifolds.
@@ 1. Introduction
[…]
and $(X_{i},A_{i})$ is a topological pair for each $1\leq i\leq m$, and combinatorial properties of $K$ is one of the key problems considered in toric topology [7]. In the particularly important and well-studied cases of a moment-angle complex $\mathcal{Z}_{K}=(\mathbb{D}^{2},\mathbb{S}^{1})^{K}$ and its real analogue $\mathcal{R}_{K}=(\mathbb{D}^{1},\mathbb{S}^{0})^{K}$, see [7, Chapter 4], it turns out that these polyhedral products are not merely cellular spaces, but also acquire a structure of a topological, or even a smooth manifold consistent with the coordinate torus action, provided $K$ is a sphere, see Definition 2.3 and Theorem 4.7.
Moreover, if the Buchstaber number $s(K)$ of a starshaped sphere $K$ (i.e. the maximal rank of a toric subgroup $H$ in the coordinate torus $\mathbb{T}^{m}$ acting freely on $\mathcal{Z}_{K}$) is maximal possible (i.e. equals mnm-n, where $n=\dim K+1$), then the topological toric manifold $\mathcal{Z}_{K}/H$ is defined. This smooth $2n$-dimensional manifold with an $n$-dimensional compact torus $T^{m}/H$ action is a topological counterpart of a toric manifold: a complete nonsingular toric variety $X_{\Sigma}$ of complex dimension $n$, which is defined in the framework of toric geometry [12] in terms of a complete regular fan $\Sigma$ in ${\mathbb{R}}^{n}$, see Definition 2.8. The real analogues of the Buchstaber invariant and the (topological) toric manifold are defined similarly.
@@
By definition, first introduced in the unpublished note [4], given an arbitrary simplicial complex $K\neq\Delta_{[m]}$ on $[m]$ with $m\geq 2$, its Bier sphere $\mathrm{Bier}(K)$ equals the deleted join of the complex $K$ and its Alexander dual complex $K^{\vee}$, see Definition 2.4. Different proofs of the fact that this construction yields a PL-sphere of dimension (m2)(m-2) appeared in [26] and [25]. More recently, a canonical starshaped realization for any Bier sphere was obtained in [19], see also [20] and Theorem 2.6. On the other hand, it was observed in [26] that almost all Bier spheres of simplicial complexes on $[m]$ are non-polytopal, as $m\to\infty$; however, no particular example of a non-polytopal Bier sphere has been constructed so far.
@@
This paper is an expansion of a short expository note [22] in which some of the present paper results, namely, Theorem 3.3, Proposition 4.3, and Proposition 4.5, were announced without proofs. We begin with a canonical regular realization for any Bier sphere and then discuss some applications of our construction in toric topology, toric geometry and geometrical combinatorics. In particular, this allows us to generalize (refine) the results on the canonical starshaped realization [19, 20] and the Buchstaber numbers computation for Bier spheres [21, 23].
The structure of the paper is as follows. In Section 2, we collect all the necessary definitions, constructions, and results related to simplicial complexes and fans. Based on that, we introduce the notion of a Bier sphere and recall its basic combinatorial and geometric properties. The central result in Section 3 is Theorem 3.3. Here we focus on the regularity of the canonical, complete fan $\Sigma_{K}$ (the radial fan of the starshaped realization of $\mathrm{Bier}(K)$, described in Theorem 2.6). As a first application we observe that whenever $\mathrm{Bier}(K)$ is polytopal the combinatorial simple polytope $P$ such that $\mathrm{Bier}(K)\cong\partial P^{*}$ is Delzant. In Section 4, we introduce the canonical (real) moment-angle manifolds and (real) toric manifolds corresponding to the canonical fan $\Sigma_{K}$ and classify all (real) canonical moment-angle manifolds of Lusternik-Schnirelmann category not greater than two (Theorem 4.9). As the first consequence of our main construction, we obtain a topological proof of the Dehn-Sommerville relations for Bier spheres based on Poincaré duality for toric manifolds (Proposition 4.3). Then an infinite family of toric manifolds that are not quasitoric is identified taking into account the canonical toric manifolds over non-polytopal Bier spheres (Proposition 4.5). Finally, a criterion for orientability of the canonical real toric manifolds over Bier spheres is proved based on the canonical characteristic matrix $\Lambda_{K}$ construction (Proposition 4.12).
@@ Acknowledgements
The authors are grateful to Anton Ayzenberg, Li Cai, Taras Panov, and Aleš Vavpetič for various fruitful discussions, valuable comments and suggestions. Limonchenko and Živaljević were supported by the Serbian Ministry of Science, Technological Development and Innovation through the Mathematical Institute of the Serbian Academy of Sciences and Arts.
@@ 2. Abstract simplicial complexes and rational polyhedral fans
Set $m]:=\{1,2,\ldots,m\}$ for an integer $m\geq 1$. We say that a non-empty subset $K\subseteq 2^{[m]}$ is an (abstract) simplicial complex on $[m]$ if $\sigma\in K$ and $\tau\subseteq\sigma$ imply $\tau\in K$. Elements of $K$ are called its simplices, or faces. Faces of cardinality $1$ are called (geometric) vertices. The set of vertices of $K$ will be denoted by $V(K)$.
@@
Given two simplicial complexes, $K_{1}$ and $K_{2}$, on $[m]$, we say that they are isomorphic and write $K_{1}\cong K_{2}$, if there exists a bijection $\phi\colon[m]\to[m]$ such that both $\phi$ and its inverse $\phi^{-1}$ map simplices to simplices. For an (n1)(n-1)-dimensional simplicial complex $K$, we define its $f$-vector to be the $(n+1)$-tuple $(f_{-1},f_{0},\ldots,f_{n-1})$ such that $f_{i}$ equals the number of $i$-dimensional faces of $K$ for $i\geq-1$. Note that $f_{-1}=1$. Then the $h$-vector of $K$ is defined to be the $(n+1)$-tuple $(h_{0},h_{1},\ldots,h_{n})$ determined from the identity:
@@
We call a pure (n1)(n-1)-dimensional simplicial complex $K$
@@
Now let us introduce the main object of study in this paper, a Bier sphere. Consider a copy $m^{\prime}]:=\{1^{\prime},2^{\prime},\ldots,m^{\prime}\}$ of the set $[m]$ with $m\geq 2$.
@@
  • the Alexander dual of $K$ is a simplicial complex $K^{\vee}$ on $m^{\prime}]$ such that
@@
  • the Bier sphere of $K$ is a simplicial complex $\mathrm{Bier}(K)$ on the disjoint union $m]\sqcup[m^{\prime}]$ such that
@@
Due to [21, Corollary 2.7], the Bier sphere $\mathrm{Bier}(K)$ has $m-f_{0}(K)+f_{m-2}(K)$ ghost vertices. Hence it has $f_{0}(\mathrm{Bier}(K))=2m-(m-f_{0}(K)+f_{m-2}(K))=m+f_{0}(K)-f_{m-2}(K)$ geometric vertices.
@@
It follows easily from the above definition that in dimension 1 the only Bier spheres are the boundaries of a $p$-gon with $p=3,4,5$, and $6$, see [24, Example 2.8]. In dimension 2, there exist $13$ different combinatorial types of Bier spheres, see [21, Theorem 2.16].
It was shown in [4] that $\mathrm{Bier}(K)$ defined above is a PL-sphere of dimension (m2)(m-2) with the number of vertices varying between $m$ and $2m$. An explicit construction of the required PL-homeomorhism between $\mathrm{Bier}(K)$ and $\partial\Delta^{m-1}$ can be found in [25]. Here is a generalization of this result, which is crucial for this paper.
@@ 3. Regular realizations for Bier spheres
Note that the statement is obvious for $m=2$. Therefore, assume $m\geq 3$. First, let us show that $\Sigma_{K}$ is a fan. By definition, $C(I,J)$ is a rational polyhedral cone generated by the set $G(I,J)$ of vectors in ${\mathbb{R}}^{m-1}$. Moreover, $G(I,J)$ is a minimal set of generators for $C(I,J)$, since $G(I,J)$ is a linearly independent set of vectors in ${\mathbb{R}}^{m-1}$.
@@
where $C^{\circ}(I,J)$ is the ”interior” of $C(I,J)\in\Sigma_{K}$, consisting of linear combinations of $e_{i}$ and $e_{j}$ with strictly positive scalars. This would also imply that the resulting fan $\Sigma_{K}$ is complete.
Here we use a relation between Bier spheres, which are sufficiently close. It is well known ([26], proof of Theorem 5.6.2) that if $K$ and $K\cup\{A\}$ are simplicial complexes with vertices in $[m]$, then:
$\mathrm{Bier}(K\cup A)=\mathrm{Bier}(K)\setminus D_{1}\cup D_{2},$
@@
    $\Sigma_{\varnothing_{[m]}}$ satisfies (3.1) (which is shown in Example 3.2, because $\Sigma_{\varnothing_{[m]}}$ is a complete regular fan);
@@
(3.2) $x=\alpha_{i_{1}}e_{i_{1}}+\cdots+\alpha_{i_{k}}e_{i_{k}}+\beta_{j_{1}}e_{j_{1}}+\cdots+\beta_{j_{l}}e_{j_{l}},$
@@
Without loss of generality, we may assume that
@@
and here $\alpha_{i_{1}}-\beta_{j_{1}}\leq 0,\ldots\alpha_{i_{k}}-\beta_{j_{1}}\leq 0,-\beta_{j_{1}}\leq 0$ and $(\beta_{j_{2}}-\beta_{j_{1}})\geq 0,\beta_{j_{l}}-\beta_{j_{1}}\geq 0$ implying that $C(A,A^{c}\setminus\{j_{1}\})$ is the cone of $\mathbb{D}_{2}$ containing $x$. Because of the uniqueness of (3.3), the point $x$ will be contained in two cones $C(A,A^{c}\setminus\{j_{1}\})$ and $C(A,A^{c}\setminus\{j_{2}\})$ if and only if $\beta_{j_{1}}=\beta_{j_{2}}$ so by disregarding the coefficients which are equal to zero in (3.3), we obtain the unique cone of $\Sigma_{K\cup\{A\}}$ containing $x$ in its interior.
@@
Since $\beta_{j_{1}}\leq 0,\alpha_{i_{2}}-\beta_{j_{1}}\leq 0,\ldots,\alpha_{i_{k-1}}-\beta_{j_{1}}\leq 0,\alpha_{m}-\beta_{j_{1}}\leq 0$ and $(\beta_{j_{2}}-\beta_{j_{1}})\geq 0,\ldots,\beta_{j_{l}}-\beta_{j_{1}}\geq 0$, we conclude that $C(A,A^{c}\setminus\{j_{1}\})$ is the cone of $\mathbb{D}_{2}$ containing $x$. The uniqueness of the cone of $\Sigma_{K\cup\{A\}}$ containing $x$ in its interior follows as in the previous case.
@@
proving that the cone $C(A,A^{c}\setminus\{m\})$, and therefore its corresponding sub-cone, contains $x$ in its interior.
@@
Recall that a (combinatorial) simple polytope is a Delzant polytope if it has a realization as a lattice polytope whose normal fan is regular. Several important classes of (convex) simple polytopes, including nestohedra [32] and 2-truncated cubes [9], were proven to consist of Delzant polytopes. As an immediate corollary of the previous theorem, we obtain the following statement.
@@ 4. Applications
Observe that a nonsingular projective toric variety $X_{\Sigma}$ corresponding to the normal fan $\Sigma=\Sigma_{P}$ of an $n$-dimensional Delzant polytope $P$ is a quasitoric manifold of real dimension $2n$ over the combinatorial polytope $P$ via restricting the action of $(\mathbb{C}^{*})^{n}$ to $\mathbb{T}^{n}\subset(\mathbb{C}^{*})^{n}$. In what follows, we consider (topological) toric manifolds with this restricted action and we denote them by $M_{\Sigma}$. The existence of the class of quasitoric manifolds arising as toric manifolds over Delzant polytopes motivated the problem, see [6, Problem 6.26], of constructing a toric manifold, which is not a quasitoric manifold. In [30], it was done by a consecutive resolution of singularities for a toric variety over the Barnette nonpolytopal 3-sphere with 8 vertices. The canonical complete regular fan $\Sigma_{K}$ construction yields a different approach to this problem.
@@
There are infinitely many canonical toric manifolds whose orbit spaces are not homeomorphic to any simple polytope as manifolds with corners.
@@
  • the smooth (m1)(m-1)-dimensional canonical real moment-angle manifold $\mathcal{R}_{\mathrm{Bier}(K)}$.
@@
In the above definition, we make use of Theorem 4.7 when $m\geq 4$. If $m\leq 3$, then $\mathrm{Bier}(K)$ is a polytopal sphere, and therefore its (real) moment-angle complex acquires an equivariant smooth structure due to [8].
Since $K$ is a full subcomplex in $\mathrm{Bier}(K)$, the previous definition and [7, Theorem 4.5.8] yield a smooth canonical moment-angle manifold with torsion in integer cohomology whenever $H^{*}(K)$ contains torsion. Similar statement holds for the existence of a nontrivial (indecomposable) triple and higher order Massey products in cohomology of such a manifold, see also [16].
Next, we are going to classify real and complex canonical moment-angle manifolds of low Lusternik-Schnirelmann category. Recall that the LS category $\mathrm{cat}(X)$ of a topological space $X$ equals the minimal integer $n$ such that $X$ can be covered by $(n+1)$ open sets $U_{i}$ contractible to a point in $X$ for each $0\leq i\leq n$. The following result shows that [23, Conjecture 1.3] holds for Bier spheres. Here we use the notation $\mathrm{cup}(X)$ for the cohomology length of a topological space $X$.
Theorem 4.9.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$.
  • (a)
    The following conditions are equivalent:
    • (1)
      $\mathrm{cat}(\mathcal{R}_{\mathrm{Bier}(K)})=1$;
    • (2)
      $\mathrm{cat}(\mathcal{Z}_{\mathrm{Bier}(K)})=1$;
    • (3)
      $\mathrm{cup}(\mathcal{R}_{\mathrm{Bier}(K)})=1$;
    • (4)
      $\mathrm{cup}(\mathcal{Z}_{\mathrm{Bier}(K)})=1$;
    • (5)
      $\mathcal{R}_{\mathrm{Bier}(K)}$ is homeomorphic to a sphere;
    • (6)
      $\mathcal{Z}_{\mathrm{Bier}(K)}$ is homeomorphic to a sphere;
    • (7)
      either $K=\varnothing_{[m]}$, or $K=\partial\Delta_{[m]}$.
  • (b)
    The following conditions are equivalent:
    • (1)
      $\mathrm{cat}(\mathcal{R}_{\mathrm{Bier}(K)})=2$;
    • (2)
      $\mathrm{cat}(\mathcal{Z}_{\mathrm{Bier}(K)})=2$;
    • (3)
      $\mathrm{cup}(\mathcal{R}_{\mathrm{Bier}(K)})=2$;
    • (4)
      $\mathrm{cup}(\mathcal{Z}_{\mathrm{Bier}(K)})=2$;
    • (5)
      $\mathcal{R}_{\mathrm{Bier}(K)}$ is homeomorphic to a connected sum of sphere products with two spheres in each product;
    • (6)
      $\mathcal{Z}_{\mathrm{Bier}(K)}$ is homeomorphic to a connected sum of sphere products with two spheres in each product;
    • (7)
      either $K$, or $K^{\vee}$ has dimension 0.
Proof.
Statement (a) follows from the fact that a closed orientable (real) moment-angle manifold over a starshaped sphere $\mathcal{K}$ has cohomology length greater than or equal to $2$ due to Poincaré duality, unless the manifold is homeomorphic to a sphere. The latter is the case if and only if $\mathrm{Bier}(K)$ is a sphere if and only if either $K$, or $K^{\vee}$ equals $\partial\Delta_{[m]}$.
[…]
@@
[…]
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Define the canonical (topological) toric manifold over $K$ and its real counterpart as a quotient space of the canonical (real) moment-angle manifold over the simplicial complex $K$:
@@
Due to [27, Theorem 1.7], the small cover $M_{{\mathbb{R}}}(P,\bar{\Lambda}_{K})$ is orientable if and only if the image of the composition $\epsilon\bar{\Lambda}_{K}$ equals $\{1\}$. This implies the statement, since $\pm e_{m}=\mp(e_{1}+\cdots+e_{m-1})\in{\mathbb{R}}^{m-1}$. If $\Sigma_{K}$ is not a normal fan of a Delzant polytope, a straightforward modification of the proof of [27, Theorem 1.7] with the starshaped region $P$ bounded by the canonical starshaped realization of $\mathrm{Bier}(K)$ constructed in [19] shows the conclusion of the orientability criterion is still valid in our case, which finishes the proof. ∎